# A convex polyhedron without Rupert’s property

### v1

@@ 1.3 Outline of the proof
Acknowledgments: We thank Alin Bostan for his everlasting moral support. Moreover, the second author is also grateful to Matthias Auchmann and A&R TECH for financial support as well as some computing resources.
@@ 3 Bounding $R_{a}(\alpha),R(\alpha),M(\theta,\varphi)$
$\|M(\theta,\varphi)P\|>\|M(\overline{\theta},\overline{\varphi})P\|-\|M(\theta,\varphi)-M(\overline{\theta},\overline{\varphi})\|\|P\|,$
@@ 4 The global theorem
$\langle R(\overline{\alpha})M(\overline{\theta}_{1},\overline{\varphi}_{2})S,w\rangle>\max_{P\in\operatorname{\mathbf{P}}}\langle M(\overline{\theta}_{2},\overline{\varphi}_{2})P,w\rangle+(\sqrt{2}+\sqrt{5})\rho\varepsilon$
@@
Consider the Octahedron $\operatorname{\mathbf{O}}=\{(\pm 1,0,0),(0,\pm 1,0),(0,0,\pm 1)\}\subseteq\operatorname{\mathbb{R}}^{3}$ and two projection directions $(\overline{\theta}_{1},\overline{\varphi}_{1})=(0,0)$ and $(\overline{\theta}_{2},\overline{\varphi}_{2})=(\pi/4,\tan^{-1}(\sqrt{2}))$. We wish to apply 17 to show that with $\overline{\alpha}=0$ this choice of parameters does not yield a solution to Rupert’s problem for $\operatorname{\mathbf{O}}$ and, moreover, that there exists $\varepsilon>0$ such that there is also no solution $(\theta_{1},\varphi_{1},\theta_{2},\varphi_{2},\alpha)$ with $|\overline{\theta}_{i}-\theta_{i}|,|\overline{\varphi}_{i}-\varphi_{i}|,|\overline{\alpha}|\leq\varepsilon$ either. In other words, the red projection in Figure 3 is not inside the black one, even after perturbing the projection parameters by some small $\varepsilon>0$.
@@ 5.1 Definitions and lemmas
For $A,\overline{A},B,\overline{B}\in\operatorname{\mathbb{R}}^{n\times n}$ and $P_{1},P_{2}\in\operatorname{\mathbb{R}}^{n}$ it holds that
@@
For $A,B\in\operatorname{\mathbb{R}}^{n\times n}$ and $P_{1},P_{2}\in\operatorname{\mathbb{R}}^{n}$ one has
@@ 5.1.2 Locally maximally distant points
Let $\varepsilon>0$ and $\theta,\overline{\theta},\varphi,\overline{\varphi}\in\operatorname{\mathbb{R}}$ with $|\theta-\overline{\theta}|,|\varphi-\overline{\varphi}|\leq\varepsilon$. Define $M=M(\theta,\varphi)$ and $\overline{M}=M(\overline{\theta},\overline{\varphi})$ and let $P,Q\in\operatorname{\mathbb{R}}^{3}$ with $\|P\|,\|Q\|\leq 1$. Then:
@@
We will show that the numerator of the left-hand side is bigger than the one of the right-hand side, and that the factors in the denominator are smaller.
@@
and observe the claim for the numerator using the triangle inequality.
@@ 5.2 Statement and proof of the local theorem
By 30 it holds that $R(\alpha)M(\theta_{1},\varphi_{2})P_{i},M(\theta_{2},\varphi_{2})Q_{i}\in\operatorname{\mathrm{Disc}}_{i}$. Moreover, 33 implies that
@@ 6.1 Definitions and lemmas
$\|R_{\operatorname{\mathbb{Q}}}(\alpha)\|,\|R^{\prime}_{\operatorname{\mathbb{Q}}}(\alpha)\|,\|X_{\operatorname{\mathbb{Q}}}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}^{\theta}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}^{\varphi}(\theta,\varphi)\|\leq 1+\kappa$
@@ 6.3 Rational Local Theorem
Then it holds that $A\geq A_{\operatorname{\mathbb{Q}}}$.
@@
It follows that the numerator of $A_{\operatorname{\mathbb{Q}}}$ is at most the numerator of $A$.
@@ 7.1 Structure of the solution tree
90 ID T #\#C ID1C sp $\overline{\theta}_{1,{\min}}$ $\overline{\theta}_{1,\max}$ $\overline{\varphi}_{1,\min}$ ${\overline{\varphi}}_{1,\max}$ ${\overline{\theta}}_{2,\min}$ ${\overline{\theta}}_{2,\max}$ ${\overline{\varphi}}_{2,\min}$ ${\overline{\varphi}}_{2,\max}$ ${\overline{\alpha}}_{\min}$ ${\overline{\alpha}}_{\max}$ $P_{1}$ $P_{2}$ $P_{3}$ $Q_{1}$ $Q_{2}$ $Q_{3}$ $r$ $\sigma_{Q}$ $w_{x}$ $w_{y}$ $w_{d}$ $S$ 0 3 4 1 1 0 64..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 1 3 30 5 2 0 16..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 2 3 30 46..67 2 16..00 32..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 3 3 30 94..77 2 32..00 48..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 4 3 30 14..51 2 48..00 64..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 5 3 4 35 3 0 16..00 0 16..00 0 64..00 0 24..00 -24..00 24..00 6 3 4 70..31 3 0 16..00 16..00 32..00 0 64..00 0 24..00 -24..00 24..00 7 3 4 10..67 3 0 16..00 32..00 48..00 0 64..00 0 24..00 -24..00 24..00 91 1 0 80..00 0 80..00 80..00 16..00 80..60 16..20 -23..40 -22..80 53..73 15..64 16..45 39 92 1 0 80..00 80..00 16..00 0 80..00 0 80..60 -24..00 -23..40 98..92 35..15 10..33 37 245 2 0 20..00 0 20..00 0 20..00 0 20..40 -22..20 -22..80 30 31 38 79 80 87 -69 1 246 1 0 20..00 0 20..00 0 20..00 20..40 40..80 -23..60 -22..20 71..05 19..88 20..37 39 247 1 0 20..00 0 20..00 0 20..00 20..40 40..80 -22..20 -22..80 71..05 19..88 20..37 39 248 2 0 20..00 0 20..00 20..00 40..00 0 20..40 -23..60 -22..20 30 31 38 79 80 87 -69 1 249 2 0 20..00 0 20..00 20..00 40..00 0 20..40 -22..20 -22..80 30 31 38 79 80 87 -69 1 18..44 1 48..00 64..00 46..00 48..00 48..00 64..00 22..80 24..00 22..80 24..00 33..40 -14..51 14..49 78

### v2

@@ 1.3 Outline of the proof
Acknowledgments: We thank Alin Bostan for his everlasting moral support. We also thank David Renshaw and Jason Reed for their immense work on the formalization of this paper in Lean and the correction of several inaccuracies of the first version of this manuscript. Moreover, the second author is also grateful to Matthias Auchmann and A&R TECH for financial support as well as some computing resources.
@@ 3 Bounding $R_{a}(\alpha),R(\alpha),M(\theta,\varphi)$
$\|M(\theta,\varphi)P\|\geq\|M(\overline{\theta},\overline{\varphi})P\|-\|M(\theta,\varphi)-M(\overline{\theta},\overline{\varphi})\|\|P\|,$
@@ 4 The global theorem
$\langle R(\overline{\alpha})M(\overline{\theta}_{1},\overline{\varphi}_{1})S,w\rangle>\max_{P\in\operatorname{\mathbf{P}}}\langle M(\overline{\theta}_{2},\overline{\varphi}_{2})P,w\rangle+(\sqrt{2}+\sqrt{5})\rho\varepsilon$
@@
Consider the Octahedron $\operatorname{\mathbf{O}}=\{(\pm 1,0,0),(0,\pm 1,0),(0,0,\pm 1)\}\subseteq\operatorname{\mathbb{R}}^{3}$ and two projection directions $(\overline{\theta}_{1},\overline{\varphi}_{1})=(0,0)$ and $(\overline{\theta}_{2},\overline{\varphi}_{2})=(\pi/4,\tan^{-1}(\sqrt{2}))$. We wish to apply 17 to show that with $\overline{\alpha}=0$ this choice of parameters does not yield a solution to Rupert’s problem for $\operatorname{\mathbf{O}}$ and, moreover, that there exists $\varepsilon>0$ such that there is also no solution $(\theta_{1},\varphi_{1},\theta_{2},\varphi_{2},\alpha)$ with $|\overline{\theta}_{i}-\theta_{i}|,|\overline{\varphi}_{i}-\varphi_{i}|,|\alpha|\leq\varepsilon$ either. In other words, the red projection in Figure 3 is not inside the black one, even after perturbing the projection parameters by some small $\varepsilon>0$.
@@ 5.1 Definitions and lemmas
For $A,\overline{A},B,\overline{B}\in\operatorname{\mathbb{R}}^{m\times n}$ and $P_{1},P_{2}\in\operatorname{\mathbb{R}}^{n}$ it holds that
@@
For $A,B\in\operatorname{\mathbb{R}}^{m\times n}$ and $P_{1},P_{2}\in\operatorname{\mathbb{R}}^{n}$ one has
@@ 5.1.2 Locally maximally distant points
Let $\varepsilon>0$ and $\theta,\overline{\theta},\varphi,\overline{\varphi}\in\operatorname{\mathbb{R}}$ with $|\theta-\overline{\theta}|,|\varphi-\overline{\varphi}|\leq\varepsilon$. Define $M=M(\theta,\varphi)$ and $\overline{M}=M(\overline{\theta},\overline{\varphi})$ and let $P,Q\in\operatorname{\mathbb{R}}^{3}$ with $\|P\|,\|Q\|\leq 1$. Assume [^1] In a first version of this manuscript we were lacking this assumption and we thank David Renshaw for pointing it out. that
$\frac{\langle\overline{M}P,\overline{M}(P-Q)\rangle-2\varepsilon\|P-Q\|\cdot(\sqrt{2}+\varepsilon)}{\big(\|\overline{M}P\|+\sqrt{2}\varepsilon\big)\cdot\big(\|\overline{M}(P-Q)\|+2\sqrt{2}\varepsilon\big)}>0,$
then:
@@
We will show that the numerator of the left-hand side is bigger than the one of the right-hand side, and that the factors in the denominator are smaller. We also show that the left-hand side is well-defined as $MP,M(P-Q)\neq 0$.
@@
and observe the claim for the numerator using the triangle inequality. In particular, $\langle{M}P,{M}(P-Q)\rangle>0$, thus $MP$ and M(PQ)M(P-Q) are non-zero, and the denominator does not vanish.
@@ 5.2 Statement and proof of the local theorem
By 30 it holds that $R(\alpha)M(\theta_{1},\varphi_{2})P_{i},M(\theta_{2},\varphi_{2})Q_{i}\in\operatorname{\mathrm{Disc}}_{i}$. Moreover, as the right-hand side in (Bε) is positive, 33 implies that
@@ 6.1 Definitions and lemmas
$\|R_{\operatorname{\mathbb{Q}}}(\alpha)\|,\|R^{\prime}_{\operatorname{\mathbb{Q}}}(\alpha)\|,\|X_{\operatorname{\mathbb{Q}}}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}^{\theta}(\theta,\varphi)\|,\|M_{\operatorname{\mathbb{Q}}}^{\varphi}(\theta,\varphi)\|\leq 1+\kappa.$
@@ 6.3 Rational Local Theorem
Assuming that $A_{\operatorname{\mathbb{Q}}}>0$, it then holds that $A\geq A_{\operatorname{\mathbb{Q}}}$.
@@
It follows that the numerator of $A_{\operatorname{\mathbb{Q}}}$ is at most the numerator of $A$ and that both are positive.
@@ 7.1 Structure of the solution tree
90 ID T #\#C ID1C sp $\overline{\theta}_{1,{\min}}$ $\overline{\theta}_{1,\max}$ $\overline{\varphi}_{1,\min}$ ${\overline{\varphi}}_{1,\max}$ ${\overline{\theta}}_{2,\min}$ ${\overline{\theta}}_{2,\max}$ ${\overline{\varphi}}_{2,\min}$ ${\overline{\varphi}}_{2,\max}$ ${\overline{\alpha}}_{\min}$ ${\overline{\alpha}}_{\max}$ $P_{1}$ $P_{2}$ $P_{3}$ $Q_{1}$ $Q_{2}$ $Q_{3}$ $r$ $\sigma_{Q}$ $w_{x}$ $w_{y}$ $w_{d}$ $S$ 0 3 4 1 1 0 64..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 1 3 30 5 2 0 16..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 2 3 30 46..67 2 16..00 32..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 3 3 30 94..77 2 32..00 48..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 4 3 30 14..51 2 48..00 64..00 0 48..00 0 64..00 0 24..00 -24..00 24..00 5 3 4 35 3 0 16..00 0 16..00 0 64..00 0 24..00 -24..00 24..00 6 3 4 70..31 3 0 16..00 16..00 32..00 0 64..00 0 24..00 -24..00 24..00 7 3 4 10..67 3 0 16..00 32..00 48..00 0 64..00 0 24..00 -24..00 24..00 91 1 0 80..00 0 80..00 80..00 16..00 80..60 16..20 -23..40 -22..80 53..73 15..64 16..45 39 92 1 0 80..00 80..00 16..00 0 80..00 0 80..60 -24..00 -23..40 98..92 35..15 10..33 37 245 2 0 20..00 0 20..00 0 20..00 0 20..40 -22..20 -22..80 30 31 38 79 80 87 955 1 246 1 0 20..00 0 20..00 0 20..00 20..40 40..80 -23..60 -22..20 71..05 19..88 20..37 39 247 1 0 20..00 0 20..00 0 20..00 20..40 40..80 -22..20 -22..80 71..05 19..88 20..37 39 248 2 0 20..00 0 20..00 20..00 40..00 0 20..40 -23..60 -22..20 30 31 38 79 80 87 955 1 249 2 0 20..00 0 20..00 20..00 40..00 0 20..40 -22..20 -22..80 30 31 38 79 80 87 955 1 18..44 1 48..00 64..00 46..00 48..00 48..00 64..00 22..80 24..00 22..80 24..00 33..40 -14..51 14..49 78